# LLM Alignment as Retriever Optimization: An Information Retrieval Perspective

### v1

@@ 8 Conclusions
This paper investigates the impact of increasing the number of retrieved passages on the performance of long-context LLMs in retrieval-augmented generation (RAG) systems. Contrary to expectations, we observe that performance initially improve but then degrade as more passages are included. This phenomenon is attributed to the detrimental influence of retrieved "hard negatives". To mitigate this issue, we propose and evaluate three solutions: training-free retrieval reordering, RAG-specific implicit LLM fine-tuning, and RAG-oriented LLM fine-tuning with intermediate reasoning. A systematic analysis of the training-based methods explores the effects of data distribution, retriever for training, and training context length. Interesting future directions include exploring (automated) position optimization with more advanced retrieval ordering methods, and fine-tuning the LLMs for RAG with more fine-grained and multi-step reasoning chains.

### v2

@@ 8 Conclusions
This paper has forged a novel link between LLM alignment and IR, offering a systematic framework to enhance the LLM alignment performance. Expanding upon this basis, we introduced LarPO, a new direct preference optimization method that integrates the IR principles to significantly enhance alignment quality. The effectiveness of LarPO is strongly supported by our comprehensive experiments across widely-used benchmarks, demonstrating its potential as a significant advancement in LLM alignment. Furthermore, our IR-focused analysis highlights the crucial role of retriever optimization objectives, hard negatives, and candidate list construction in achieving effective alignment.
@@
Acknowledgements
This research was supported in part by Apple PhD Fellowship, in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, in part by the Office of Naval Research contract number N000142412612, in part by NSF grant numbers IIS-19-56151 and 2402873, in part by the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897 and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329, in part by Cisco, and in part by the Center for Intelligent Information Retrieval. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of the sponsors or the U.S. Government.